OpenStudy (anonymous):

What is the mean number of hours that high school students spend on homework per week? It is known that the standard deviation of the population is 2.2 hours. In a random sample of 40 students, it is found that they average 13 hours per week on homework. a) What is the standard deviation for this sample? b) What is the critical value for a 90% confidence interval? c) What is the critical value for a 95% confidence interval?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

a) use the formula s = sigma/sqrt(n) in this case, sigma = 2.2 and n = 40

OpenStudy (anonymous):

) 2.2/sqrt(40) = 0.3478 <--- for a

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's the std dev for the sample of 40

jimthompson5910 (jim_thompson5910):

one sec while i get a calc ready

OpenStudy (anonymous):

ok so thats a

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i seee .9?

jimthompson5910 (jim_thompson5910):

before that, to the left a bit

jimthompson5910 (jim_thompson5910):

P(|z| < ....)

jimthompson5910 (jim_thompson5910):

what goes in the .... place

jimthompson5910 (jim_thompson5910):

it's on the line where you see "confidence level"

OpenStudy (anonymous):

1.645?

jimthompson5910 (jim_thompson5910):

good, that's your critical value for the 90% CI CI = confidence interval

jimthompson5910 (jim_thompson5910):

now change that 0.90 to 0.95 and recompute

OpenStudy (anonymous):

ok so b is 1.645?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

tell me what you get for c

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

1.96

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

ook cool! d?

jimthompson5910 (jim_thompson5910):

i don't see part d

OpenStudy (anonymous):

here:

OpenStudy (anonymous):

d) What is the 90% confidence interval? Be sure to include the formula with your numbers plugged in. Use the model: x plus or minus z times sigma divided by the square root of n e) What is the 95% confidence interval?

jimthompson5910 (jim_thompson5910):

ok for part d, we'll break it into two parts

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

first part lower limit L = xbar - C*s L = 13 - 1.645*0.3478 ... note we found the value for s back in part a) L = 12.427869

jimthompson5910 (jim_thompson5910):

second part upper limit U = xbar + C*s U = 13 + 1.645*0.3478 U = 13.572131

jimthompson5910 (jim_thompson5910):

So the CI (L, U) turns into (12.427869, 13.572131) ie (L, U) = (12.427869, 13.572131)

jimthompson5910 (jim_thompson5910):

therefore, the 90% CI is (12.427869, 13.572131) and you can round this however you need to or leave it like it is

jimthompson5910 (jim_thompson5910):

for part e), you're doing the exact same thing as part d...but...you're using 1.96 for C instead of 1.645 for C basically, C = 1.96 in part e) while everything else stays the same

OpenStudy (anonymous):

so D is (12.427869, 13.572131)

jimthompson5910 (jim_thompson5910):

yeah pretty much

OpenStudy (anonymous):

ok so E?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do the same thing in part d), but use C = 1.96 this time

OpenStudy (anonymous):

whats C?

OpenStudy (anonymous):

can u write it ouut? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

just go back to the work for D and instead of using C = 1.645, use C = 1.96

OpenStudy (anonymous):

L = 13 minus 1.96 * 0.3478 = 12.427869 for the upper limit: U = 13 plus 1.96 * 0.3478 = 13.57213

OpenStudy (anonymous):

well not the equal part, but the formula?

jimthompson5910 (jim_thompson5910):

good, you just need to fix the results and you're done

OpenStudy (anonymous):

13- .681688= 12.318312

OpenStudy (anonymous):

13.681688

jimthompson5910 (jim_thompson5910):

good on both

jimthompson5910 (jim_thompson5910):

so your 95% CI is (12.318312, 13.681688)

OpenStudy (anonymous):

is it ANSWER = (12.318312, 13.681688)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

@jim_thompson5910 can u help with a fe wmore